You may need to use the appropriate technology to answer this question.

Researchers at two universities found that airlines are doing a better job of getting passengers to their destinations on time. Company A and Company B Airlines were among the leaders in on-time arrivals with both having 88% of their flights arriving on time. But for the 12% of flights that were delayed, how many minutes were these flights late? Sample data showing the number of minutes that delayed flights were late are shown below for both airlines.

Company A
34 59 43 30 3
32 42 85 30 48
110 50 10 26 70
52 83 78 27 70
27 90 38 52 76
Company B
44 65 42 34 67
104 45 27 38 84
76 44 33 51 63
42 35 34 64 65

(a)

Formulate the hypotheses that can be used to test for a difference between the population mean minutes late for delayed flights by these two airlines. (Let μ1 = population mean minutes late for delayed Company A flights and μ2 = population mean minutes late for delayed Company B flights.)

H0: μ1μ2 ≤ 0

Ha: μ1μ2 > 0

H0: μ1μ2 < 0

Ha: μ1μ2 = 0

    

H0: μ1μ2 ≥ 0

Ha: μ1μ2 < 0

H0: μ1μ2 = 0

Ha: μ1μ2 ≠ 0

H0: μ1μ2 ≠ 0

Ha: μ1μ2 = 0

(b)

What is the sample mean number of minutes late for delayed flights for each of these two airlines?

Company A minCompany B min

(c)

Calculate the test statistic. (Round your answer to three decimal places.)

What is the p-value? (Round your answer to four decimal places.)

p-value =

Using a 0.05 level of significance, what is your conclusion?

Reject H0. There is statistical evidence that one airline does better than the other in terms of their population mean delay time.Do not reject H0. There is no statistical evidence that one airline does better than the other in terms of their population mean delay time.    Reject H0. There is no statistical evidence that one airline does better than the other in terms of their population mean delay time.Do not Reject H0. There is statistical evidence that one airline does better than the other in terms of their population mean delay time.
Answer #1

For Company A, sample 1 :  
x̅1 =   50.6
s1=   26.56596
n1 =   25
For Company B, sample 2 :  
x̅2 =   52.85
s2 =   20.01914
n2 =    20
α =    0.05

a) Null and Alternative hypothesis:
Ho : µ1 - µ2 = 0
Ha : µ1 - µ2 ≠ 0

b) x̅1 - x̅2 = 50.6 -52.85 = -2.25

c) Pooled variance :  
S²p = ((n1-1)*s1² + (n2-1)*s2² )/(n1+n2-2)  
=   570.9897

Test statistic:              
t = (x̅1 - x̅2)/[√(s²p(1/n1 + 1/n2 )]              
= -0.314

df = n1+n2-2 = 43   
              
p-value :              
At t = -0.314 and df = 43, two tailed p-value =T.DIST.2T(-0.3139, 43) = 0.7551

Conclusion:

Do not reject H0. There is no statistical evidence that one airline does better than the other in terms of their population mean delay time.